At the instant, the traffic light turns green, a truck with a constant velocity of 12 m/s. At the same moment, the car begins to accelerate at 1 m/s^2. when it reaches 20 m/s, the car continues at this velocity. When does the car pass the truck? How far have they gone from the traffic light? 


1
Expert's answer
2021-09-29T09:54:05-0400

Given:

"v_i=12\;\rm m/s"

"v_f=20\;\rm m/s"

"a=1\;\rm m/s^2"


The time of motion

"t=\frac{v_f-v_i}{a}=\frac{20-12}{1}=8\:\rm s"

The distance traveled by a car

"d=\frac{v_f^2-v_i^2}{2a}=\frac{20^2-12^2}{2*1}=128\:\rm m"